QUESTION 8 $500 450 E 400 350 250 2 200 쵸 E 150 C 100 U 50 45° $50 100 150 200 250 300 350 400 450 500 550 Real Domestic Output ($B) Refer to the graph for a private closed economy. At the equilibrium level of GDP, saving will be O $100 billion. $50 billion. O $150 billion. an amount that cannot be determined from the information given
Answer #1

Ans: $100 billion

Explanation:

In a closed economy

When RGDP is $0, C = $50 billion. It means savings = -$50 billion (it is called dissaving)

At equilibrium

C + Ig = C + S

450 = 350 + S

S = 450 - 350 = $100 billion